OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. k^2-k-2/ k^2-4k-5

OpenStudy (anonymous):

k-2/k-5 < answer but what are the restricted numbers ?

OpenStudy (anonymous):

@agent0smith

OpenStudy (aum):

k cannot have values that would make the denominator, k^2-4k-5, zero because division by zero is not allowed.

OpenStudy (agent0smith):

Brackets, again :P always around the entire denominator and entire numerator: (k^2-k-2)/ (k^2-4k-5)\[\large \frac{k^2-k-2} {k^2-4k-5}\]now we gotta factor everything

OpenStudy (anonymous):

yeah i know my answer is \[\frac{ k-2 }{ k-5 }\]

OpenStudy (anonymous):

but idk what the resricted number are

OpenStudy (anonymous):

its either -1 and 5 or -1 and -5

OpenStudy (agent0smith):

Oh okay. What value of k makes the denominator equal to zero? The denominator is k-5...

OpenStudy (anonymous):

5 ?

OpenStudy (aum):

What values of k in the original problem, before simplification, that would make the denominator zero?

OpenStudy (aum):

Denominator = k^2-4k-5 = (k-5)(k+1) Denominator will be zero when k = 5 or -1 So k cannot be 5 or -1.

OpenStudy (agent0smith):

Yes, you need the original factored form, before simplifying, to get all the restricted values. k+1 cancels off, but k can't equal -1